OpenStudy (joftheworld):

An archaeologist in Turkey discovers a spear head that contains 32% of it's original amount of C-14. Find the age of the spear head to the nearest year

OpenStudy (aaronq):

for a first order reaction, you can use the exponential growth/decay equation: \(A_{t}=A_0*e^{-kt}\) where \(A_{t}\) is the amount after t time elapsed, \(A_0\) is the initial amount. t is time, and k is the decay constant (see below). -------------------------------------------------- Alternatively, you can use the first-order integrated rate law: \(\large ln[A]_t=-kt+ln[A]_o\) where \(ln[A]_o\) is the initial concentration of the substance \(k\) is the decay constant \(t\) is time \(ln[A]_t\) is the concentration of substance at time \(t\) ------------------------------------------- \(\large t_{1/2}=\dfrac{ln(2)}{k}\). \(t_{1/2}\) is the half-life

OpenStudy (aaronq):

great, so you have all the variables. plug them into the equation and solve it

OpenStudy (aaronq):

Or you can rearrange them first, then plug your values in. Either works

OpenStudy (anonymous):

\[0.32=e^{-.0001t}\] solve for \(t\)

OpenStudy (joftheworld):

0.0001 * t = in(1/32) ?

OpenStudy (aaronq):

yep it's ln not in

OpenStudy (joftheworld):

so t = (1/32)/0.0001 ?

OpenStudy (aaronq):

dont forget the ln \(t = \dfrac{ln(1/32)}{0.0001}\)

OpenStudy (aaronq):

oh ps, it's 0.32 not 32 \(t = \dfrac{ln(1/.32)}{0.0001}\)

OpenStudy (joftheworld):

i got it now :D

OpenStudy (aaronq):

dope